OpenStudy (anonymous):

CAN SOMEONE CHECK IF I HAVE THE RIGHT ANSWER? for Find the derivative of the function. y=Ine^x

zepdrix (zepdrix):

\[y=\ln(e^x)\]like that? :o

OpenStudy (anonymous):

y(prime)=1/xe^x+e^x1/x

OpenStudy (anonymous):

no w.out the ( )

zepdrix (zepdrix):

is the power inside the log or no? :)

OpenStudy (anonymous):

nope :l as it is written initially

OpenStudy (anonymous):

y=Ine^x

zepdrix (zepdrix):

are you allowed to do any simplifying before you take the derivative? or is derivation suppose to be your first step? because there is a really nice simplification here :o

OpenStudy (anonymous):

umm u simplify 1st

OpenStudy (amistre64):

spose a = e^x a' = e^x D [ln(a)] = a' /a

zepdrix (zepdrix):

\[\ln e=1\] gives you a nice place to start :D y=1^x

OpenStudy (amistre64):

not 1^x

OpenStudy (amistre64):

if anything ln(e^x) = x ln(e) = x

zepdrix (zepdrix):

his equation is y=(ln e)^x

OpenStudy (amistre64):

really? hmm

OpenStudy (anonymous):

wont i 1st rewrite In?

OpenStudy (amistre64):

ln e then is just a constant ...

OpenStudy (anonymous):

???????????????????

zepdrix (zepdrix):

lol that one always confuses people :)

zepdrix (zepdrix):

you need to get comfortable with natural log lilly!! :D ln e, means --> e to what power = e

OpenStudy (amistre64):

should go something like this eh \[a^x=y\] \[x=log_a(y)\] \[x=\frac{ln(y)}{ln(a)}\] \[x~ln(a)=ln(y)\]---------- derive \[ln(a)=\frac{y'}{y}\] \[y~ln(a)=y'\] \[a^x~ln(a)=y'\]

OpenStudy (anonymous):

okay now i KNOW i am failing that quiz on FRIDAY ......:[[[[[[[[[[

OpenStudy (amistre64):

interactmath.com its a free site similar to MathLab. lots and lots of practice problems there to work on :)

OpenStudy (anonymous):

THANKS ILL CHECK IT OUT :]

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

okay no i do need help

OpenStudy (amistre64):

that site walks you thru the problems as well, giving step by step solutions.

OpenStudy (anonymous):

my calc book is not on the site

OpenStudy (amistre64):

it doesnt have to be, the material in a calculus book hasnt changed for cenuries

OpenStudy (amistre64):

solving derivatives in one calculus books is the same as solving them in all the others

OpenStudy (anonymous):

whta book do u suggest i click on

OpenStudy (amistre64):

let me give it a look see.

OpenStudy (amistre64):

http://interactmath.com/ChapterContents.aspx adams calculus is the first title i see that has to do with calculus, open up to view the contents

OpenStudy (amistre64):

i see bittinger calculus and applications down the list

OpenStudy (anonymous):

and what ch do u suggest i click on?

OpenStudy (anonymous):

on adams

OpenStudy (amistre64):

briggs cochran as well yeah, try out the adams book

OpenStudy (amistre64):

heres a screenshot of what im looking at

OpenStudy (amistre64):

most of the calculus titles look about the same; and there are about a dozen that i spotted in the list

OpenStudy (anonymous):

plz help

OpenStudy (amistre64):

plz help doing what?

OpenStudy (anonymous):

i cant click on any of the things 4 the website it wont let me

OpenStudy (amistre64):

i wouldnt know how to help with that ... it works fine for me

OpenStudy (anonymous):

okay thnx

OpenStudy (amistre64):

if youve got a textbook, try working thru the problems at the end of the chapters, and not just the assigned ones

OpenStudy (amistre64):

post them up if youve got questions or ideas about them

OpenStudy (anonymous):

okay :)